### v1

# The discovery and characterization of Earth-crossing asteroid 2024 YR4

Abstract
We describe observations and physical characteristics of Earth-crossing asteroid 2024 YR4, discovered on 2024 December 27 by the Asteroid Terrestrial-impact Last Alert System. The asteroid has semi-major axis, = 2.52 au, eccentricity, = 0.66, inclination = 3.41∘, and a 0.003 au Earth minimum orbit intersection distance. We obtained g, r, i, and Z imaging with the Gemini South/Gemini Multi-Object Spectrograph on 2025 February 7. We measured a g-i spectral slope of 133 /100 nm, and color indices g-r = 0.700.10, r-i = 0.250.06, i-Z = -0.270.10. 2024 YR4 has a spectrum that best matches R-type and Sa-type asteroids and a diameter of 30-65 m using our measured absolute magnitude of 23.90.3 mag, and assuming an albedo of 0.15-0.4. The lightcurve of 2024 YR4 shows 0.4 mag variations with a rotation period of 1170 s. We use photometry of 2024 YR4 from Gemini and other sources taken between 2024 December to 2025 February to determine the asteroid’s spin vector and shape, finding that it has an oblate, 3:1 a:c axial ratio and a pole direction of , = 42∘, -25∘. Finally, we compare the orbital elements of 2024 YR4 with the NEO population model and find that its most likely sources are resonances between the inner and central Main Belt.
1 Introduction
Earth-crossing asteroids (ECAs) are asteroids whose orbital trajectories cross the Earth’s coming within its 0.983 au perihelion, , and 1.017 au aphelion, , distances. By definition, ECAs belong to the Apollo NEO class with semi-major axes, 1 au, and 1.017 au, and the Aten NEO class with 1 au and 0.983 au. ECAs present scientific opportunities for detailed study of close approaching near-Earth objects (NEO, e.g., Granvik et al., 2013; Binzel et al., 2015). Interactions with the Earth’s tides during close-approaching asteroids can affect the structural integrity of asteroids, causing changes in their surface material properties (Binzel et al., 2010), or completely disrupt, making Earth-encounters a source of some NEOs (Granvik & Walsh, 2024; Nesvorný et al., 2024). Additionally, a portion of EC asteroids are a source of impacts on the Earth and the Moon affecting their total impact flux (Mazrouei et al., 2019; Cohen et al., 2023). Thus, ECAs provide examples of useful study on the effect of close approaches on the physical properties of asteroids and the impact chronology of the Earth and Moon.
The majority of ECAs originate from sources in the inner Main Belt between 2.2 au and 2.5 au (Granvik et al., 2018). Asteroids leave the Main Belt due to the Yarkovsky effect causing their semi-major axes to evolve, increasing for prograde-spinning asteroids and decreasing for asteroids that are spinning retrograde (Vokrouhlický et al., 2015). As asteroids evolve in their semi-major axis due to the Yarkosvky effect, they cross resonances such as the resonance that defines the inner border of the Main Belt at 2.2 au, the 3:1 mean motion resonance (MMR) with Jupiter at 2.5 au, and various weaker resonances located between 2.2 au and 2.5 au (Granvik et al., 2017). The interaction between the orbits of these asteroids and the resonances causes their orbits to chaotically evolve, sending them on trajectories where they can encounter the Earth (Wisdom, 1983; Farinella et al., 1993).
While the majority of inner Main Belt asteroids belong to the siliceous S-type group, 20 of inner Main Belt asteroids belong to the more primitive and carbonaceous C-type group DeMeo & Carry (2013). This results in the majority of C-type NEOs originating from the inner region of the Main Belt (Binzel et al., 2019). Additionally, a considerable portion of NEOs can also belong to the X-type class, whose compositions span enstatite achondrite and metallic compositions (Bus & Binzel, 2002; Avdellidou et al., 2022). While the orbits of NEOs and their likely sources within the Main Belt provide constraints on their composition (Jedicke et al., 2022), the information on the composition of NEOs can be enhanced through observations covering visible and near-infrared wavelengths (DeMeo et al., 2009).
This paper discusses the discovery of ECA 2024 YR4 by the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al., 2018) telescope and observations taken at Gemini South and the determination of its physical shape model. We also adapt the techniques described by Bolin et al. (2020, 2021, 2025) of using multi-band photometry to constrain the physical and taxonomic properties of 2024 YR4. Finally, we will use the lightcurve inversion techniques established by Hanuš et al. (2011, 2018) on publically sparse and dense data to determine the spin vector and shape of 2024 YR4, and test its likely source within the Main Belt using the NEO population model (Morbidelli et al., 2020; Nesvorný et al., 2023, 2024).
2 Observations
The ATLAS Chile, Rio Hurtado telescope, with Minor planet Center (MPC) observing code W68, made the first discovery observations of 2024 YR4 on 2024 December 27 05:42:49 UTC as seen in Panels a-d of Fig 1. The asteroid was discovered at right ascension (RA) and declination (dec) of 08:56:40.97, -00:16:11.93, and was located 0.996 au from the Sun, 0.017 au from the Earth, and at a phase angle of 43.2∘, and an apparent magnitude of 16.17 in the ATLAS o-band filter (Denneau et al., 2024). The ATLAS camera has a pixel scale of 1.86 ″, an “orange” o-band custom filter providing wavelength coverage between 560 nm and 820 nm, and an effective wavelength of 663 nm (Tonry et al., 2018). The asteroid was detected in all four o-band images taken in a 3160 s interval from 2024 December 27 05:42:49 UTC to 2024 December 27 06:35:30 UTC. The asteroid was moving northwest at 22″/min, resulting in the detections forming a trail 11″ long (panels a-d, Fig. 1).
A director’s discretionary time request (DDT) was granted at Gemini South to observe 2024 YR4 on 2025 February 7 02:23:46 UTC under program GS-2025A-DD-102 (PI: Bolin). The asteroid was located at RA, dec = 08:04:08.66, +13:33:24.31, 1.367 au from the Sun, 0.399 au from the Earth and had a phase angle of 14.8∘ during the 2025 February 7 UTC observations. The Gemini Multi-Object Spectrograph (GMOS) instrument (Hook et al., 2004) was used in imaging mode to observe 2024 YR4 in g, r, I, and Z filters. The GMOS camera comprises a 5.5′ 5.5′ field of view Hamamatsu array with an effective pixel scale of 0.08″. The camera was used in 2 2 binning mode, giving an effective pixel scale of 0.16″. The GMOS g, r, and i filers are equivalent to Sloan Digital Sky Survey (SDSS) g, r, and i filters with effective wavelengths of 475 nm, 630 nm, and 780 nm (Fukugita et al., 1996). A filter was used for coverage past 800 nm equivalent to the the WFCAM Z with an effective wavelength of 880 nm (Casali et al., 2007). This filter has the advantage of being equivalent to the SDSS z filter, but avoids telluric absorption features near 950 nm (Hodgkin et al., 2009). Observations of 2024 YR4 were also taken on 2025 February 23 02:13:17 UTC when the asteroid was located at RA, dec = 08:04:08.66, 13:33:24.31, 1.520 au from the Sun, 0.598 au from the Earth, and had a phase angle of 21.7∘, but only exposures in i-band were taken. During both the 2025 February 7 UTC and 2025 February 23 UTC observing dates, 200 s exposures were used for all filters, and the telescope was tracked at the rate of motion of 2024 YR4.
For this manuscript, we will focus mainly on the characterization of 2024 YR4 for which we obtained observations with Gemini South on 2025 February 7 UTC. During the observations, the measured seeing in the r band was 0.8″ and the airmass ranged from 1.43 -1.50 during the 2 h 35 min observing sequence. We acquired 11 g band images, 12 r band images, 8 i band images, and 8 Z band images. The g, r, i, and Z filters changed between exposures in a rgiZrgriZgr sequence to smear out the brightness variations in the asteroid’s lightcurve and limit its effect on the color measurements. The DRAGONS image pipeline was used with bias and flat field calibration frames to detrend the data (Labrie et al., 2023). Images of 2024 YR4 when it intersected a background star were discarded. A mosaic showing 2024 YR4 in the co-added g, r, i, and Z frames is shown in Fig 2.
3 Results
3.1 Orbit and source population
Astrometry of 2024 YR4 from observing dates was reported to the MPC (Williams, 2025a, b). Additional astrometry was reported by 440 observations of 2024 YR4 between 2024 December 25 UTC and 2025 March 1 UTC 111https://minorplanetcenter.net/db_search/show_object?utf8=%E2%9C%93&object_id=2024+YR4, accessed on 2025 March 4 UTC.. Table 1 shows the six orbital element solution computed by JPL HORIZONS using these observations222https://ssd.jpl.nasa.gov/tools/sbdb_lookup.html#/?sstr=2024%20YR4, solution date 2025 March 1 UTC, accessed on 2025 March 4 UTC.. The asteroid has an Apollo-type orbit with 1 au, and a perihelion takes it inside the orbit of the Earth. The Earth minimum orbit intersection distance (MOID) according to JPL HORIZONS is 2.82 au.
| Heliocentric Elements | |
| Epoch (JD) | 2,460,800.5 |
| Time of perihelion, (JD) | 2,460,636.9175586(3.76x10-5) |
| Semi-major axis, (au) | 2.5158656(2.35x10-5) |
| Eccentricity, | 0.66154790(3.35x10-6) |
| Perihelion, (au) | 0.851499970(4.87x10-7) |
| Aphelion, (au) | 4.1802311(3.90x10-5) |
| Inclination, (∘) | 3.4081731(1.22x10-5) |
| Ascending node, (∘) | 271.36561725(8.89x10-6) |
| Argument of perihelion, (∘) | 134.3613663(1.54x10-5) |
| Mean Anomaly, (∘) | 40.402628(5.7x10-4) |
The orbital and location of 2024 YR4 at the time of its discovery on 2024 December 27 UTC is shown on Fig. 3 based on values from JPL Horizons333https://ssd.jpl.nasa.gov/horizons/app.html#/. We compared the orbit elements, , , , and absolute magnitude (see following Section 3.2) of 2024 YR4 with the NEOMOD3 population model (Nesvorný et al., 2024), suggesting its most likely source is the 3:1 MMR with Jupiter resonance located in between the inner and central Main Belt regions at 2.5 au with a 73.9 probability. The next most likely location, with a 9.8 chance is the resonance located at the inner edge of the Main Belt at 2.2 au. We combine our source probabilities for 2024 YR4 with the NEO albedo model of Morbidelli et al. (2020) to estimate its albedo, , finding a predicted value of 0.18.
3.2 Color photometry and spectral classification
We measured the photometry of 2024 YR4 in the Gemini g, r, i, and Z images using a 1.12″ aperture and a sky-median subtraction annulus with and inner and outer radius of 2.08″ and 2.88″. Local solar analog stars from the Pan-STARRS catalog (Chambers et al., 2016) were used to calibrate the photometry according to the transformations described by Tonry et al. (2012). The brightnesses per bandpass of our measured photometry are g = 24.010.09, r = 23.310.04, i = 23.060.04, Z = 23.340.09. The asteroid’s g-i color and spectral gradient are g-i = 0.950.10 and 133/100 nm. The other multi-band color indices of 2024 YR4 are g-r = 0.700.10, r-i = 0.250.06, i-Z = -0.270.10, r-Z = -0.020.10, and the a∗ parameter, an indicator of spectral slope defined by Ivezić et al. (2001) as a∗ = (0.89 (g-r)) + (0.45 (r-i)) - 0.57 is a∗ = 0.160.09.
The a∗ and i-z(Z) of 2024 YR4 are plotted in Fig. 4, which has a∗ and i-z(Z) on the border between S-types, which have on average a∗ = 0.120.03 and i-z = -0.08 0.07, and V-types, which have on average a∗ = 0.15 0.11 and i-z = -0.46 0.04 (DeMeo et al., 2009). However, this is still redder compared to the Sun which has a∗ = -0.13 and i-z = 0.03 (Willmer, 2018), and C-types, which have on average a∗ = -0.090.02 and i-z = 0.020.03 (DeMeo et al., 2009). Additionally, the equivalent V band brightness of 2024 YR4 is V = 23.920.10 determined by combining our g and r band measurements of the asteroid and the color transformation from (Jordi et al., 2006).
We estimate the H magnitude of 2024 YR4 using our estimated magnitude and the HG phase function (Bowell et al., 1988):
where is the 1.367 au heliocentric distance of 2024 YR4, is its geocentric distance of 0.399 au and is its phase angle of 14.8∘ of 2024 YR4 on 2025 February 7 UTC. is the phase coefficient which we use the value of 0.23, the average value for S-type asteroids (Pravec et al., 2012; Vereš et al., 2015). and are the basis functions normalized at = 0∘ described in (Bowell et al., 1988). We obtain = 23.910.10 in agreement from the value calculated from JPL HORIZONS of = 23.960.28 but caution that the uncertainty on estimate of is underestimated due to the lack of information about the phase function (e.g., Bolin et al., 2022). Therefore, we adopt the larger uncertainty on taken from JPL HORIZONS of 0.3.
For each g, r, i, and Z filter, we compute the spectral reflectance of 2024 YR4 , dividing its measured flux by the flux of a solar analog star. The reflectivity spectrum of 2024 YR4 normalized to a wavelength of 550 nm is shown in Fig. 5, and features a red slope from 470 nm to 770 nm, and absorption at 880 nm. We take the approach of Bolin et al. (2023) by comparing the spectrum of 2024 YR4 with the Bus-DeMeo asteroid spectral catalog (DeMeo et al., 2009) as a test of its taxonomic spectral type. We determine the reduced statistic between the reflectance of 2024 YR4 and S-type asteroids (S, Sq, Sv, Q), C-types (B, C, Cg, Cgh), X-types (X, Xc, Xe, Xk, Xn), and assorted other asteroid types (A, D, K, L, O, R, V). The closest match with the spectrum of 2024 YR4 is to the R-type and Sa-type spectral classes with a reduced of 0.9, 1.1. The next closest match after R-type and Sa-type asteroids is with Sr-type asteroids with a reduced of 1.7. By comparison, the reduced for the spectrum of 2024 YR4 with V-types and C-types are 2.5 and 11.8.
As a rough estimate of 2024 YR4’s size, we estimate its visible albedo, , assuming it has similar albedo properties as other small S-complex and R-type NEOs of 0.15-0.40 (Delbo et al., 2017). This range contains the independent albedo estimate for 2024 YR4 that we obtained by using with the NEO albedo model (Morbidelli et al., 2020) of 0.18 when considering the 0.1 spread in albedos of small NEOs (Delbo et al., 2003). We use the albedo estimate and H magnitude of 2024 YR4 to determine the diameter, D, of 2024 YR4 using the equation from Harris & Lagerros (2002) arriving at D = 30-65 m.
3.3 Lightcurve photometry and shape model
The g, r, i, and Z GMOS photometry data of 2024 YR4 from 2025 February 7 UTC were used to search for periodic variations in its brightness. The colors computed from Section 3.2 were used to convert the g, i, and Z data into equivalent r-band magnitudes. The combined r-band equivalent g, r, i, and Z time series lightcurve are plotted in the top panel of Fig. 6. The lightcurve shows evidence of variations larger than the 0.1 mag 1- photometric uncertainties of the individual data points in the lightcurve.
We take a similar approach as described by Bolin et al. (2024) to estimating the lightcurve period of the r-band equivalent data by applying a Lomb-Scargle periodogram (LS, Lomb, 1976; Scargle, 1982) algorithm and a phase dispersion minimization algorithm (PDM, Stellingwerf, 1978) to the data. The LS and PDM algorithms result in finding an estimated lightcurve period of 587-590s as seen in the top and bottom panels of Fig. 7. We use the approach of (Purdum et al., 2021) estimate the uncertainty on the determined period by removing data points from the time series lightcurve and repeating our periodogram estimation of the lightcurve period 1,000 times, resulting in a central value of 586 s and a 1 uncertainty estimate of 142 s. These period estimates imply that 2024 YR4 has a double-peaked synodic rotation period of 1172284 s.
The central panel of Fig. 6 shows the r-band equivalent lightcurve of 2024 YR4 folded by the double-peaked rotation period of 1172 s. The result shows almost complete coverage in the lightcurve of 2024 YR4 over its 1172 s rotation. We rebin the phased lightcurve of 2024 YR4 into 0.075 wide bins in phase dimensional space as seen in the bottom panel of Fig. 6. We note that estimating an asteroid’s shape from a single lightcurve observation is affected by the degeneracy between viewing angle and axial ratio (e.g., Bolin et al., 2018b), therefore, we will take a multi-epoch approach to estimating the asteroid’s shape as described below.
textbfLeft panel: Lomb-Scargle periodogram of lightcurve period vs. spectral power (Lomb, 1976) for the g filter data from the 2024 September 27 UTC Gemini N/GMOS observations. A peak in power is located at a single-peaked lightcurve period of 587 s. Right panel: Phase dispersion minimization analysis of lightcurve rotation period vs. metric (Stellingwerf, 1978). The metric is minimized at single-peaked rotation periods of 1150 s, similar to the 1320227 s period found with the Lomb-Scargle Periodogram.
To estimate a more accurate shape for 2024 YR4, we use the lightcurve inversion method (Kaasalainen et al., 2001; Kaasalainen & Torppa, 2001), which has been widely applied to derive the physical properties of small bodies, including near-Earth asteroids (NEAs). A similar approach was successfully used in the case of (3200) Phaethon, where photometric data from multiple observing geometries determined its rotational state and shape (Hanuš et al., 2018a).
For 2024 YR4, we used a combination of dense lightcurves obtained from Gemini (MPC observatory code, I11) we used only data in the r and g filters that are the most accurate), public data taken by the Very Large Telescope (MPC observatory code 309), and sparse photometry reported to the Minor Planet Center (MPC) that provided sufficient coverage of observing geometries. The modeling procedure followed the methodology outlined in our previous studies (Athanasopoulos et al., 2022; Hanuš et al., 2023).
We used photometric observations reported to the MPC: only data from observatories providing magnitude measurements with at least two decimal places were considered to ensure the highest possible accuracy. After initial modeling attempts, we further filtered out the MPC photometry, ending with three data sets. Namely, observations from multiple ATLAS (Asteroid Terrestrial-impact Last Alert System; Tonry et al., 2018) stations (MPC observatory codes (T05, T08, and W68) were combined into a single data set in the o filter, consisting of 25 individual measurements. Additional observations from the Catalina Sky Survey (Larson et al., 2003) (MPC observatory code 703), Pan-STARRS2 (MPC observatory code F52) (Chambers et al., 2016) stations combined together in the G filter finalized the sparse dataset.
The photometric data were processed following the methodology outlined in Hanuš et al. (2011) to prepare them for lightcurve inversion. This procedure included transformation to fluxes, epoch light-travel time correction, flux normalization to unit distance to the Sun and the observer, and unity. Any outliers or measurements affected by systematic errors were identified and removed.
Dense photometric observations of 2024 YR4 were obtained from the European Southern Observatory (ESO) Science Archive as part of a long-term monitoring program (ESO Program ID 113.2690.002, PI Hainaut). The program, conducted in Service Mode at VLT UT1-Antu, supports the ESA Space Situational Awareness Near-Earth Objects Protection Program, which aims to characterize the physical properties of potentially hazardous asteroids. The observations were taken with the FORS2 instrument on 2025 January 21 UTC.
The VLT data are publicly available, and we retrieved them from the ESO database as raw images. The data were processed using the MPO Canopus software package, following standard photometric reduction procedures (Warner et al., 2009). Bias and flat-field corrections were applied to all frames, followed by aperture photometry extraction, calibrated against known field stars to ensure accurate magnitude measurements.
The available photometric dataset spans 42 days of observations, covering a range of phase angles from 45.7∘ to 14.8∘, which is critical for constraining the shape model. Table A.1 summarizes the complete list of photometric measurements, including observation epochs, filter bands, and sources.
We conducted a broad period search within the range of 0.15 to 0.50 hours, encompassing the suspected synodic period of approximately 0.32 hours derived from the photometric dataset. A well-defined minimum was identified in the periodogram at hours (1168s) (Figure 8). This solution remains stable against variations in the relative weighting of the individual datasets, confirming its robustness.
Adopting this sidereal period, we explored possible spin-vector orientations by testing a wide range of initial values isotropically distributed on a sphere. The results indicate a degree of sensitivity to the chosen data weighting. Several pole solutions were tested using different relative weight configurations, with the most prominent solution found at an ecliptic longitude and latitude of approximately and , respectively. This particular solution satisfies physical constraints, ensuring that the derived shape model rotates around its principal axis with maximum momentum of inertia. While alternative pole solutions cannot be entirely excluded, this solution appears to be the most plausible. This solution will be included in the Database of Asteroid Models from Inversion Techniques (DAMIT) lightcurve inversion software package (Durech et al., 2015).
Figure 9 presents the fit to the photometric data, incorporating dense lightcurves obtained from VLT and Gemini in the r and g filters, along with sparse photometric measurements from ATLAS, the Catalina Sky Survey, and Pan-STARRS. The derived shape of 2024 YR4 is notably peculiar, resembling a highly flattened spheroid with nearly equal equatorial dimensions (Fig. 8).
4 Discussion and Conclusions
The properties of 2024 YR4 are similar to other small NEOs, that it has a S-complex composition, most likely originated from a resonances located inner Main Belt, and likely has moderate reflective properties implying a size of 42 m. The exact origin within the Main Belt seems less clear, though a clue may lie in its spin-vector orientation. As recent asteroid lightcurve spin-vector studies have shown, the spin direction of an asteroid can affect the way an asteroid’s orbit evolves due to the thermal recoil Yarkovsky effect, with prograde asteroids drifting outward, and retrograde asteroids drifting inwards (Bolin et al., 2018a; Hanuš et al., 2018b; Athanasopoulos et al., 2022). Since the spin direction of 2024 YR4 is retrograde with its -39∘ ecliptic latitude, an origin from the inner Main Belt since its retrograde spin would cause it to drift inwards away from the 3:1 MMR at 2.5 au, its most likely escape path into the NEO population. Therefore, its suspected retrograde spin may imply that it original location was the central Main Belt, located at 2.52 au to 2.82 au from the Sun (Nesvorný et al., 2015). This would take 2024 YR4 ’s origins away from the inner Main Belt where S-types dominate to the central Main Belt where C-types are more plentiful seemingly in tension with its taxonomic type (DeMeo & Carry, 2014). However, several S-type families are located near the 3:1 MMR in the central Main Belt, such as the Rafita family located at 2.58 au, which could be a candidate source family for 2024 YR4.
acknowledgments
This study is based on observations obtained at the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation on behalf of the Gemini Observatory partnership.
Gemini Observatory is located on Maunakea, land of the Knaka Maoli people, and a mountain of considerable cultural, natural, and ecological significance to the indigenous Hawaiian people. The authors wish to acknowledge the importance and reverence of Maunakea and express gratitude for the opportunity to conduct observations from the mountain.
This work has been supported by the grant 25-16789S of the Czech Science Foundation.
The authors wish to thank Matthew Belyakov for help with the observaitons of 2024 YR4, and Mauro Barbieri for help with searching the VLT data archive. The authors also wish to thank K. Wierzchos, and C. Fuls for their help with measuring the astrometry from the Gemini images and submitting it to the MPC. The authors also wish to think the team members of the Catalina Sky Survey, the Panoramic Survey Telescope and Rapid Response System, an ATLAS for submitting observations of 2024 YR4 to the MPC which were used in this work.
This research has made use of data and/or services provided by the International Astronomical Union’s Minor Planet Center.
ATLAS, Gemini:South, VLT:Antu